# Beyond mirkwood: Enhancing SED Modeling with Conformal Predictions

### 21 Dec 2023 v1

@@ Discussion
Conclusions and Future Work
This study marks a substantial advancement in the field of spectral energy distribution (SED) fitting by integrating flexible machine learning models, particularly CatBoost, with the innovative technique of conformalized quantile regression. This approach not only enhances the accuracy of SED fitting but also introduces a new depth to the uncertainty quantification in astronomical research. The adaptability of our tool to various astronomical datasets, coupled with the ability to select from a range of sklearn-compatible models, ensures its applicability across different research contexts. CatBoost’s effectiveness in handling complex datasets, combined with our sophisticated method of uncertainty quantification, allows for more reliable and nuanced interpretations of galactic properties.

### 10 Feb 2024 v2

@@ 5.4 Discussion
6 Conclusions and Future Work
This study marks a substantial advancement in the field of spectral energy distribution (SED) fitting by integrating flexible machine learning models, particularly CatBoost, with the innovative technique of conformalized quantile regression. This approach not only enhances the accuracy of SED fitting but also introduces a new depth to the uncertainty quantification in astronomical research. The adaptability of our tool to various astronomical datasets, coupled with the ability to select from a range of scikit-learn-compatible models, ensures its applicability across different research contexts. CatBoost’s effectiveness in handling complex datasets, combined with our sophisticated method of uncertainty quantification, allows for more reliable and nuanced interpretations of galactic properties.
@@ 6 Conclusions and Future Work
[…]
This research received no external funding.
[…]
Not applicable.
[…]
Not applicable.
[…]
The raw data supporting the conclusions of this article will be made available by the authors on request.
[…]
The author declares no conflicts of interest.
[…]